OpenStudy (anonymous):

sarah spends $28 on food, this amount is 4/9 of his allowance. calculate his allowance

OpenStudy (stormfire1):

\[\frac{4}{9}x=28\]\[\frac{9}{4}28=x=$63\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you help in another question?

OpenStudy (stormfire1):

ok

OpenStudy (anonymous):

a model of a car has a scale of 1:25. the model is 18cm long. calculate in metres the actual length of the car

OpenStudy (stormfire1):

If the model is 18cm and is 1/25th of size of the actual car, the actual car length is \[18cm*25=450cm\]That's more commonly shown as 4.5m.

OpenStudy (stormfire1):

In other words, if the scale is 1:25 then for every cm of model, there are 25 cm of actual car. Make sense?

OpenStudy (anonymous):

yeah, i guess so

OpenStudy (anonymous):

so its 4.5?

OpenStudy (stormfire1):

4.5m or 450cm...both are equivalent.

OpenStudy (anonymous):

oh ok, 1 more

OpenStudy (stormfire1):

ok

OpenStudy (anonymous):

camilla has $5 to spend in the market. she buys 1 and half kg of bananas priced at 80 cents per kg and 3 yams priced at 45 cents each. how much money does she have?

OpenStudy (stormfire1):

Well let's see. For the bananas:\[\frac{0.80c}{kg}*1.5kg=$1.20\]For the yams: \[0.45c * 3 = $1.35\]\[$5-$1.35-$1.20=$2.45\]So she has $2.45 left.

OpenStudy (anonymous):

i didnt get the banana part :S

OpenStudy (anonymous):

you there?

OpenStudy (stormfire1):

I was answering another question. I'm back :)

OpenStudy (anonymous):

ok :)

OpenStudy (stormfire1):

ok...let's see. What part didn't you understand about the bananas?

OpenStudy (anonymous):

why we divided 0.80 over kg

OpenStudy (stormfire1):

That's just another way of writing the fact that the bananas are 0.80 center PER kg...the PER implies division.

OpenStudy (stormfire1):

err *cents

OpenStudy (anonymous):

oh ok

OpenStudy (stormfire1):

We're not actually dividing the amount by "kg"...it's just a way of expressing the ratio. It also helps to keep the units straight later on when you take higher courses.

OpenStudy (anonymous):

a block of cheese, of mass 8 kilograms, is cut by machine into 500 equal slices. calculate the mass of one slice of cheese in kg

OpenStudy (stormfire1):

Ok, you just need to divide the total mass by 500. What do you get?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

500 divide 8?

OpenStudy (stormfire1):

The block of cheese is 8kg...that's your total mass. You want to know the mass of one slice after you divide that block into 500 pieces. Mathematically, just divide 8 by 500.

OpenStudy (anonymous):

i got it :)

OpenStudy (stormfire1):

\[\frac{8kg}{500}\]

OpenStudy (stormfire1):

ok :)

OpenStudy (anonymous):

write in its simplest form, the ratio 3.5kg: 800 grams

OpenStudy (anonymous):

we have to convert right?

OpenStudy (stormfire1):

Ok, now we're back to ratios. Remember when I said a ratio can be written as a fraction? Well, you can do the same thing here but first you need to convert one of the units to match the other. You either convert all the grams to kilograms or vice versa. I'll convert the grams to kilograms as follows:\[1kg = 1000g so 3.5kg = 3.5*1000=3500g\] Now that out units are right, write the ratio as a fraction:\[\frac{3500g}{800g}=\frac{4.375}{1}\]The one on the right is in simplest form.

OpenStudy (anonymous):

its 4.375:1?

OpenStudy (stormfire1):

Yep :)

OpenStudy (anonymous):

my math teacher's gna be proud :P

OpenStudy (stormfire1):

What grade? Just curious

OpenStudy (anonymous):

8th grade lol :{P

OpenStudy (stormfire1):

Well, trust me on this: Pay attention in math, science and english...those will get you the furthest in life :) The rest of the class....blah.

OpenStudy (anonymous):

but math is REALLY hard, i got a 4.4/10 on my test, and most of the questions i got right were from my friend :/

OpenStudy (stormfire1):

Maybe you should get a tutor and practice more then. Math is the most useful class of all....but that's just my opinion.

OpenStudy (anonymous):

yeah, i guess so :/

OpenStudy (stormfire1):

Getting better at math really just involves doing a lot of practice problems...that's the only way to learn it. It doesn't take a genius...just practice.

OpenStudy (anonymous):

can you help me in one more question?

OpenStudy (stormfire1):

ok

OpenStudy (anonymous):

it has 2 parts

OpenStudy (stormfire1):

ok

OpenStudy (anonymous):

there are approximately 500,000 grains of wheat in 2kg of bag. a) calculate the mass of 1 grain in grams

OpenStudy (stormfire1):

That's just like the problem we did a moment ago. You have the total mass (2kg) and the number of pieces to break it into (500,000). What do you think you need to do?

OpenStudy (anonymous):

divide?

OpenStudy (stormfire1):

Yes....go on

OpenStudy (anonymous):

2/500,000?

OpenStudy (stormfire1):

Yes!

OpenStudy (anonymous):

:D :D

OpenStudy (stormfire1):

What did you get?

OpenStudy (anonymous):

4*10-6

OpenStudy (stormfire1):

That's correct...and you should know that it should be a very small value because you broke the 2kg into 500,000 pieces.

OpenStudy (anonymous):

how do we right it not in standard form?

OpenStudy (anonymous):

whole number

OpenStudy (stormfire1):

Since it's 10^-6 you have to move the decimal place 6 places to the left: 0.000004kg

OpenStudy (anonymous):

but they said we have to write it in grams

OpenStudy (stormfire1):

No problem...how many grams are in a kilogram?

OpenStudy (anonymous):

1000?

OpenStudy (stormfire1):

Yes...so multiply your answer times 1000.

OpenStudy (anonymous):

0.000004*1000?

OpenStudy (stormfire1):

yes

OpenStudy (anonymous):

4*10-3 is 0.0004?

OpenStudy (anonymous):

0.004?

OpenStudy (stormfire1):

That's better :)

OpenStudy (stormfire1):

So it's 0.004 grams

OpenStudy (anonymous):

:) thank you, i have to go do the rest of my homework :)

OpenStudy (stormfire1):

Ok, good luck!

OpenStudy (anonymous):

thank you, bye